OpenStudy (gabylovesyou):

A cell phone plan offers a monthly rate of $25 for unlimited minutes plus $.05 per text message. Which algebraic expression below represents the total cost for this cell phone plan in terms of "x"?

OpenStudy (gabylovesyou):

A. 25.05x B. 25x +0.05 C. 25 + 0.05x D. 24.95x

OpenStudy (gabylovesyou):

@asnaseer @AccessDenied @AnimalAin @agentc0re help plis

jimthompson5910 (jim_thompson5910):

let x = number of text messages sent

jimthompson5910 (jim_thompson5910):

so if you send 1 message, then it will cost 0.05 dollars plus the monthly rate of $25 So one message will cost 0.05 + 25 = 25.05 dollars --------------- 2 messages will cost 0.05+0.05 or 0.05*2 = 0.1 dollars Add this onto 25 to get 25+0.1 = 25.10 dollars total --------------- 3 messages will cost 0.05+0.05+0.05 or 0.05*3 = 0.15 dollars Add this onto 25 to get 25+0.15 = 25.15 dollars total --------------- See the pattern? In general, sending x text messages will cost 0.05x dollars which adds to the $25 to get 0.05x+25 dollars total

jimthompson5910 (jim_thompson5910):

Of course adding can be done in any order, so 0.05x+25 is the same as 25+0.05x

OpenStudy (gabylovesyou):

ohhh..... ok!! so its C :)

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (gabylovesyou):

Thank you!!! :)

jimthompson5910 (jim_thompson5910):

you can confirm this answer by plugging in (or replacing x with) various numbers

jimthompson5910 (jim_thompson5910):

and getting the corresponding costs

OpenStudy (gabylovesyou):

ok!! :) thanks

jimthompson5910 (jim_thompson5910):

np